Can you help me solve this Linear Combination Equation?
-2x+3y=-9
x+3y=3

Answer by ewatrrr(24785)

 
Hi,
-2x+3y= -9
x+3y= 3 |Subtracting 2nd from 1st EQ to eliminate y
-3x = -12
x = 4 and y = -1/3 (4 + 3y = 3, 3y = -1, y = -1/3)
CHECKING our answer***
Using: -2x+3y= -9
-8 + -1 = -9